# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v1

@@ 3.1 Overview
The remainder of this paper will consist of the proofs of Theorem A and Theorem B. For the main arguments, it will be convenient to work in a local coordinate system in which $\mathcal{M}$ corresponds to a linear subspace. Such a coordinate system will be introduced in section 3.2. Theorem A (i) then follows from an elementary argument presented in section 3.3. In order to prove Theorem A (ii), in section 3.4, we employ a center-stable manifold theorem. In section 3.5, we introduce a random dynamical systems framework to treat stochastic gradient descent. Theorem B (i) is then proved in section 3.6. The proof is similar to the proof of Theorem A (i) presented in section 3.3, but the possibility for non-uniform hyperbolicity adds an additional challenge. Finally, Theorem B (ii) is proved in sections 3.7-3.9. The proof is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (see e.g. [BS88, Bax91]). While these works rely on Hörmander conditions [Hör67], we use a criterion by LePage [LP82] to establish a spectral gap for the projective semi-group. The assumptions for the criterion of LePage are checked in section 3.7. In section 3.8, we construct a local Lyapunov function similar to the recent works [BBPS22, BCZG23]. The proof of Theorem B is then completed in section 3.9.
@@ 3.2 Local Coordinates
Proof.
Let $A\in\mathbb{R}^{D\times(D-N)}$ be some matrix which induces an isomorphism from $\mathbb{R}^{D-N}$ onto $\mathcal{T}(x^{*})$. Since $\mathcal{M}$ is a embedded smooth manifold, we can find neighborhoods $0\in U_{\mathcal{T}}\subseteq\mathcal{T}(x^{*})$ and $x^{*}\in\tilde{U}\subseteq\mathbb{R}^{D}$ and smooth map $\zeta:U_{\mathcal{T}}\to\mathcal{N}(x^{*})$ with $\zeta(0)=0$, $\zeta^{\prime}(0)=0$ and
@@
so in particular $\chi$ is a local diffeomorphism and (3.1) holds. Set $\hat{V}:=\chi^{-1}(\tilde{U})$ and $\hat{U}:=\chi(\hat{V})$. If we now let $\chi$ be its restriction $\chi:\hat{V}\to\hat{U}$, it is a smooth diffeomorphism, satisfying i) to iii). If $\chi$ is not bi-Lipschitz we can reassign $\hat{V}$ to a smaller neighborhood, which is pre-compact in the original $\hat{V}$. After reassigning $\hat{U}$ and restricting $\chi$ accordingly, $\chi$ will be bi-Lipschitz. ∎
@@
Proof.
It is sufficient to consider each of the functions $\hat{\mathcal{G}}_{\eta},\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ separately, find a neighborhood $V_{\delta}$ on which the corresponding inequality from (3.6) or (3.7) holds and conclude the proof by choosing the intersection over all these $V_{\delta}$. In the following, we will only consider $\hat{\mathcal{G}}_{\eta}$ and will show how to find an $V_{\delta}$ such that (3.6) holds, as the proofs for $\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ work analogously.
@@
This finishes the proof. ∎
@@ 3.3 Gradient Descent - the Stable Case
Proof of Theorem A (i).
Suppose $\mu(x^{*})<0$ and let $U\subseteq\mathcal{M}$ be some neighborhood of $x^{*}$. Our goal is to prove $\mathbb{P}(X^{\operatorname{GD}}_{\lim}\in U)>0$. Choose $\delta>0$ such that $e^{\mu(x^{*})}+\delta=:\gamma<1$, let $V_{\delta}$ be the corresponding neighborhood given by Lemma 3.2. Let $R_{\delta}>0$ be some radius, such that
@@ 3.4 Gradient Descent - the Unstable Case
In the case $\mu(x^{*})<0$ considered in the previous section, all eigenvalues of $\mathds{1}_{N}-\eta G_{x^{*}}$ lie strictly inside the unit circle. As a consequence we were able to construct an open neighborhood of $x^{*}$, such that for every initial condition in that neighborhood, GD converges to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$. In the case $\mu(x^{*})>0$ we consider in this section, we only know that some eigenvalue of $\mathds{1}_{N}-\eta G_{x^{*}}$ lies strictly outside the unit circle while there could still be eigenvalues on or strictly inside the unit circle. As a consequence on should not expect to be able to construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for any initial condition in $U$. Instead we will construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for Lebesgue-almost any initial condition in $U$. In fact, we will show that the set of initial conditions in $U$ for which GD does converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ is contained in a lower dimensional manifold called the center-stable manifold [^10] While the set, we call center-stable manifold is indeed a $C^{1}$-manifold (cf. [KH95, Theorem 6.2.8]), we will only show that it is the graph of a Lipschitz continuous function.. While for the preceding and all subsequent sections it is convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a $D-N$-dimensional tangential part $v$ and an $N$-dimensional transversal part $w$, here it will be more convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a center-stable part $v_{-}$ corresponding to the eigenvalues strictly inside or on the unit circle and an unstable part $v_{+}$ corresponding to the eigenvalues strictly outside the unit circle.
@@
Theorem 3.3 (center-stable manifold).
@@
If $\bar{\mathcal{G}}_{\eta}$ is a local diffeomorphism at 0, this result is standard (see e.g. [KH95, chapter 6.2]). In our setting, however, it is possible that $\bar{\mu}_{i}=0$ for some $i\in[D]$. Still, Theorem 3.3 follows from the arguments in [KH95] under mild modifications. Alternatively, stable/unstable/center manifold theorems in non-locally diffeomorphic setting have also been obtain by reducing the problem to the locally diffeomorphic case by an abstract method called inverse limits ([RS80, QXZ09]).
@@
Proof.
By Hypothesis (H2), we have $\mathbb{P}(X_{0}^{\operatorname{GD/SGD}}\in A)=\nu(A)=0$. By Hypothesis (H3), we also have $\mathbb{P}(X_{n}^{\operatorname{GD/SGD}}\in A)=0$, for every $n\geq 1$ and thus
@@
Proof of Theorem A (ii).
Suppose $\mu(x^{*})>0$. Let $x^{*}\in U\subseteq\mathbb{R}^{D}$ be the open neighborhood given by
@@ 3.5 Random Dynamical System Framework for SGD
While in sections 3.3 and 3.4 we studied the dynamics of gradient descent, the rest of the paper will be concerned with establishing the analogous results for stochastic gradient descent. For this, we will formulate SGD as a random dynamical system (see [Arn98] for a general introduction into the theory). Formally, for some appropriate probability space $(\Omega,\mathcal{F},\mathbb{P})$, we will introduce a map
@@
A map $\theta:\Omega\to\Omega$ satisfying (3.16) is called metric dynamical system. Equation (3.15) is known as the cocycle property and a map $\varphi$ satisfying it is called a cocycle over $\theta$. A pair $(\theta,\varphi)$, consisting of a metric dynamical system and a cocycle over it, is called random dynamical system (cf. [Arn98, Definition 1.1.1]).
@@
whenever the left-hand side is well defined, i.e. whenever $(n+m,\omega,v,w)\in D_{\psi}$. This turns the pair $(\theta,\psi)$ into a local random dynamical system (cf. [Arn98, Definition 1.2.1]). The local cocycle $\psi$ can be seen as the local lift of $\varphi$ via $\chi$, as one can easily check that
@@ 3.6 Stochastic Gradient Descent - the Stable Case
satisfying $\mathbb{P}(C_{\gamma}(\omega)<\infty)=1$. ∎
@@
Proof of Theorem B (i).
Suppose $\lambda(x^{*})<0$ and let $U\subseteq\mathcal{M}$ be some neighborhood of $x^{*}$. Our goal is to prove $\mathbb{P}(X^{\operatorname{SGD}}_{\lim}\in U)>0$. Choose $\gamma\in\mathbb{R}$, such that $e^{\lambda(x^{*})}<\gamma<1$ and let $\delta>0$ and $C_{\gamma}:\Omega\to\mathbb{R}\cup\infty$ be such that the conclusion of Lemma 3.6 holds, i.e. such that $\mathbb{P}(\mathbb{C}_{\gamma}(\omega)<\infty)=1$ and that we have
@@ 3.7 Generated Matrix Semi-groups
It remains to prove Theorem B (ii). The proof of Theorem B (i) in the previous section was of a quenched nature. One might expect that the best approach to proving Theorem B (ii) is to construct $\omega$-wise center-stable manifolds similar to the ones constructed in section 3.4. While an invariant manifold theory (Pesin theory) has been developed for random dynamical systems [LQ95], these results only provide center-stable manifolds for single points $x\in\mathcal{M}$. For the argument in section 3.4 it was crucial to have a center-stable manifold for an open subset of $\mathcal{M}$. The authors are not aware of any method to construct such a random center-stable manifold.
@@
We will show this by constructing a Lyapunov function (see e.g. [BH22]) defined on a neighborhood of $x^{*}$ which goes to infinity near $\mathcal{M}$. This is inspired by previous work on the instability of invariant subspaces for stochastic differential equations [Bax91, BS88, BBPS22, BCZG23, CZH21]. While these works rely on conditions of Hörmander-type to establish the existence of a spectral gap in the so-called projective process, such a spectral gap is excluded by the discrete nature of our problem. Instead we use an argument due to LePage [LP82] (cf. also [BL85] for a survey in English) to find a spectral gap in a different topology. LePage’s argument needs the matrix semi-group on which the linear co-cycle $\Psi$ is supported to satisfy two algebraic properties, namely being contracting and strongly irreducible (cf. Definition 3.8 below). In this section, we will show that they follow from the regularity of $x^{*}$.
@@
Proof.
Since $x^{*}\in\mathcal{M}$ is regular, in particular
@@
Thus $\mathcal{S}_{1}(x^{*})\subseteq\operatorname{GL}(N,\mathbb{R})$ and since $\mathcal{S}(x^{*})$ is generated by $\mathcal{S}_{1}(x^{*})$ also $\mathcal{S}(x^{*})\subseteq\operatorname{GL}(N,\mathbb{R})$. ∎
Definition 3.8.
@@
Proof.
Let $x^{*}\in\mathcal{M}$ be a regular point with $\lambda(x^{*})>0$. For ease of notation, we will write $G:=G_{x^{*}}$ and $G_{[i]}:=G_{x^{*},[i]}$. We start by proving that $\mathcal{S}(x^{*})$ is contracting. From the original definition (2.12) of $\lambda(x^{*})$, we get the inequality
@@
Since $G_{1,1}\neq\frac{2}{\eta}$, we have $|1-\eta G_{1,1}|\neq 1$ and the sequence $(\kappa_{n})$ consists of pairwise distinct elements. Thus, in particular, the sequence $(W_{n})$ consists of pairwise distinct subspaces, completing the proof. ∎
@@ 3.8 Construction of Lyapunov Functions
Following recent works in fluid dynamics [BBPS22, BCZG23], we will study the family of linear operators $(\mathcal{P}_{q}:\mathcal{C}^{0}(S^{N-1})\to\mathcal{C}^{0}(S^{N-1}))_{q\in\mathbb{R}}$ given by
@@
Here $\mathcal{C}^{0}(S^{N-1})$ denotes the Banach space of real-valued continuous functions on $S^{N-1}$, equipped with the supremum-norm. It can be readily seen that the operators $\mathcal{P}_{q}$ are bounded in this norm [^14] In fact $\|\mathcal{P}_{q}\|\leq K_{2}^{q}$ for $q\geq 0$ and $\|\mathcal{P}_{q}\|\leq K_{1}^{-q}$ for $q\leq 0$, with $K_{1}$ and $K_{2}$ defined by (3.28) and (3.29).. The arguments in [BBPS22, BCZG23] rely on the fact that $\mathcal{P}_{0}$ admits a spectral gap in their settings. Unfortunately the discrete nature of our setting prohibits such a spectral gap, at least in $\mathcal{C}^{0}(S^{N-1})$. In order to circumvent this obstacle we must consider a different Banach space.
@@
Lemma 3.10 ([BL85, Part A-Proposition V.4.1]).
@@
For the proof, we refer to [BL85]. The integrability assumption in [BL85, Part A-Proposition V.4.1] is satisfied trivially in our setting, as $\Psi^{(1)}_{\omega}$ can only take finitely many values.
Lemma 3.10 (ii) says that 1 an dominant eigenvalue of $\mathcal{P}_{0}$, i.e. it is isolated and the spectral value with the largest absolute value. By classical perturbation theory ([Kat95]) this implies that for $q$ sufficiently close to 0 the largest spectral value of $\mathcal{P}_{q}$ is also an isolated eigenvalue and both the dominant eigenvalue and the corresponding Riesz projections are analytic in $q$. Put more precisely, we get the following corollary (cf. also [BL85, Part A-Theorem V.4.3]). Here $\dot{\mathcal{C}}^{\alpha}$ denotes the dual space of $\mathcal{C}^{\alpha}$ and $\mathcal{Q}^{*}\in L(\dot{\mathcal{C}}^{\alpha})$ the dual operator of an operator $\mathcal{Q}\in L(\mathcal{C}^{\alpha})$.
Corollary 3.11.
@@
The arguments for the rest of the section are similar to the ones made in [BBPS22, chapter 4]. Henceforth, let $\alpha\in(0,1)$ be as in Lemma 3.10.
Theorem 3.12.
@@
Proof.
Note that (3.42) states that $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$. In the following, we will show that $\frac{\mathrm{d}}{\mathrm{d}q}r(q)|_{q=0}=\lambda(x^{*})>0$. This will imply that for sufficiently small $p>0$, we can set $\gamma:=r(-p)<1$ and $f^{*}:=f_{-p}$. Since, for small $p$, the function $f_{-p}$ is close151515in the $\mathcal{C}^{\alpha}$-sense, but thus in particular in the $\mathcal{C}^{0}$-sense to $f_{0}=\mathbf{1}$, it is indeed a positive function and by (3.41), we have $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$.
@@
By the Oseledet’s theorem ([Ose68], cf. also [Arn98, Theorem 3.4.11]) the limit in the last line is equal to $\lambda(x^{*})$. Thus we get $r(q)\geq e^{q\lambda(x^{*})}$, for sufficiently small $q$. Since we already know that $r$ is differentiable in 0 from Corollary 3.11 and since $r(0)=1$, this implies $\frac{\mathrm{d}}{\mathrm{d}q}r(q)|_{q=0}=\lambda(x^{*})$, completing the proof. ∎
@@
Proof.
This is a direct consequence of (3.42). Let $w\in\mathbb{R}^{N}\setminus\{0\}$ and set $s:=\frac{w}{\|w\|}\in S^{N-1}$. Then
@@
Proof.
Using $\alpha$-Hölder continuity of $f^{*}$ and the bounds in (3.45), we can estimate
@@
Proof.
Let $\varepsilon>0$ and choose $\tilde{\delta}>0$ such that the conclusion (3.47) of the previous lemma holds. Recall that (cf. (3.28))
@@
which implies (3.49). ∎
Corollary 3.16.
@@
Proof.
First note that if we choose $\delta<K_{1}$, using Corollary 3.5, we get
@@
for all $\omega\in\Omega$ and all ${(v,w)\in V_{\delta}\setminus(\mathbb{R}^{D-N}\times\{0\})}$. ∎

### v2

@@ 3.1 Overview
The remainder of this paper will consist of the proofs of Theorem A and Theorem B. For the main arguments, it will be convenient to work in a local coordinate system in which $\mathcal{M}$ corresponds to a linear subspace. Such a coordinate system will be introduced in section 3.2. Theorem A (i) then follows from an elementary argument presented in section 3.3. In order to prove Theorem A (ii), in section 3.4, we employ a center-stable manifold theorem. In section 3.5, we introduce a random dynamical systems framework to treat stochastic gradient descent. Theorem B (i) is then proved in section 3.6. The proof is similar to the proof of Theorem A (i) presented in section 3.3, but the possibility for non-uniform hyperbolicity adds an additional challenge. Finally, Theorem B (ii) is proved in sections 3.7-3.9. The proof is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (see e.g. Baxendale and Stroock (1988); Baxendale (1991)). While these works rely on Hörmander conditions (cf. Hörmander (1967)), we use a criterion by Le Page (1982) to establish a spectral gap for the projective semi-group. The assumptions for the criterion of Le Page are checked in section 3.7. In section 3.8, we construct a local Lyapunov function similar to the recent works by Bedrossian et al. (2022) and Blumenthal et al. (2023). The proof of Theorem B is then completed in section 3.9.
@@ 3.2 Local Coordinates
Proof  Let $A\in\mathbb{R}^{D\times(D-N)}$ be some matrix which induces an isomorphism from $\mathbb{R}^{D-N}$ onto $\mathcal{T}(x^{*})$. Since $\mathcal{M}$ is a embedded smooth manifold, we can find neighborhoods $0\in U_{\mathcal{T}}\subseteq\mathcal{T}(x^{*})$ and $x^{*}\in\tilde{U}\subseteq\mathbb{R}^{D}$ and smooth map $\zeta:U_{\mathcal{T}}\to\mathcal{N}(x^{*})$ with $\zeta(0)=0$, $\zeta^{\prime}(0)=0$ and
@@
so in particular $\chi$ is a local diffeomorphism and (3.1) holds. Set $\hat{V}:=\chi^{-1}(\tilde{U})$ and $\hat{U}:=\chi(\hat{V})$. If we now let $\chi$ be its restriction $\chi:\hat{V}\to\hat{U}$, it is a smooth diffeomorphism, satisfying i) to iii). If $\chi$ is not bi-Lipschitz we can reassign $\hat{V}$ to a smaller neighborhood, which is pre-compact in the original $\hat{V}$. After reassigning $\hat{U}$ and restricting $\chi$ accordingly, $\chi$ will be bi-Lipschitz.
@@
Proof  It is sufficient to consider each of the functions $\hat{\mathcal{G}}_{\eta},\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ separately, find a neighborhood $V_{\delta}$ on which the corresponding inequality from (3.6) or (3.7) holds and conclude the proof by choosing the intersection over all these $V_{\delta}$. In the following, we will only consider $\hat{\mathcal{G}}_{\eta}$ and will show how to find an $V_{\delta}$ such that (3.6) holds, as the proofs for $\hat{\mathcal{G}}_{\eta,1},\dots,\hat{\mathcal{G}}_{\eta,N}$ work analogously.
@@
This finishes the proof.
@@ 3.3 Gradient Descent - the Stable Case
Proof [of Theorem A (i)] Suppose $\mu(x^{*})<0$ and let $U\subseteq\mathcal{M}$ be some neighborhood of $x^{*}$. Our goal is to prove $\mathbb{P}(X^{\operatorname{GD}}_{\lim}\in U)>0$. Choose $\delta>0$ such that $e^{\mu(x^{*})}+\delta=:\gamma<1$, let $V_{\delta}$ be the corresponding neighborhood given by Lemma 3.2. Let $R_{\delta}>0$ be some radius, such that
@@ 3.4 Gradient Descent - the Unstable Case
In the case $\mu(x^{*})<0$ considered in the previous section, all eigenvalues of $\mathds{1}_{N}-\eta G_{x^{*}}$ lie strictly inside the unit circle. As a consequence we were able to construct an open neighborhood of $x^{*}$, such that for every initial condition in that neighborhood, GD converges to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$. In the case $\mu(x^{*})>0$ we consider in this section, we only know that some eigenvalue of $\mathds{1}_{N}-\eta G_{x^{*}}$ lies strictly outside the unit circle while there could still be eigenvalues on or strictly inside the unit circle. As a consequence on should not expect to be able to construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for any initial condition in $U$. Instead we will construct a neighborhood $U\subseteq\mathbb{R}^{D}$ of $x^{*}$, such that GD does not converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ for Lebesgue-almost any initial condition in $U$. In fact, we will show that the set of initial conditions in $U$ for which GD does converge to some $X_{\lim}^{\operatorname{GD}}\in\mathcal{M}$ near $x^{*}$ is contained in a lower dimensional manifold called the center-stable manifold [^11] While the set, we call center-stable manifold is indeed a $C^{1}$-manifold (cf. Theorem 6.2.8 in Katok and Hasselblatt (1995)), we will only show that it is the graph of a Lipschitz continuous function.. While for the preceding and all subsequent sections it is convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a $D-N$-dimensional tangential part $v$ and an $N$-dimensional transversal part $w$, here it will be more convenient to consider vectors in $\mathbb{R}^{D}$ as consisting of a center-stable part $v_{-}$ corresponding to the eigenvalues strictly inside or on the unit circle and an unstable part $v_{+}$ corresponding to the eigenvalues strictly outside the unit circle.
@@
Theorem 3.3 (Center-Stable Manifold)
@@
If $\bar{\mathcal{G}}_{\eta}$ is a local diffeomorphism at 0, this result is standard (see e.g.  chapter 6.2 in Katok and Hasselblatt (1995)). In our setting, however, it is possible that $\bar{\mu}_{i}=0$ for some $i\in[D]$. Still, Theorem 3.3 follows from the arguments in Katok and Hasselblatt (1995) under mild modifications. Alternatively, stable/unstable/center manifold theorems in non-locally diffeomorphic setting have also been obtain by reducing the problem to the locally diffeomorphic case by an abstract method called inverse limits (see e.g. Ruelle and Shub (1980) or Qian et al. (2009)).
@@
Proof  By Hypothesis (H2), we have $\mathbb{P}(X_{0}^{\operatorname{GD/SGD}}\in A)=\nu(A)=0$. By Hypothesis (H3), we also have $\mathbb{P}(X_{n}^{\operatorname{GD/SGD}}\in A)=0$, for every $n\geq 1$ and thus
@@
Proof [of Theorem A (ii)] Suppose $\mu(x^{*})>0$. Let $x^{*}\in U\subseteq\mathbb{R}^{D}$ be the open neighborhood given by
@@ 3.5 Random Dynamical System Framework for SGD
While in sections 3.3 and 3.4 we studied the dynamics of gradient descent, the rest of the paper will be concerned with establishing the analogous results for stochastic gradient descent. For this, we will formulate SGD as a random dynamical system (see Arnold (1998) for a general introduction into the theory). Formally, for some appropriate probability space $(\Omega,\mathcal{F},\mathbb{P})$, we will introduce a map
@@
A map $\theta:\Omega\to\Omega$ satisfying (3.16) is called metric dynamical system. Equation (3.15) is known as the cocycle property and a map $\varphi$ satisfying it is called a cocycle over $\theta$. A pair $(\theta,\varphi)$, consisting of a metric dynamical system and a cocycle over it, is called random dynamical system (cf. Defintion 1.1.1 in Arnold (1998)).
@@
whenever the left-hand side is well defined, i.e. whenever $(n+m,\omega,v,w)\in D_{\psi}$. This turns the pair $(\theta,\psi)$ into a local random dynamical system (cf. Definition 1.2.1 in Arnold (1998)). The local cocycle $\psi$ can be seen as the local lift of $\varphi$ via $\chi$, as one can easily check that
@@ 3.6 Stochastic Gradient Descent - the Stable Case
satisfying $\mathbb{P}(C_{\gamma}(\omega)<\infty)=1$.
@@
Proof [of Theorem B (i)] Suppose $\lambda(x^{*})<0$ and let $U\subseteq\mathcal{M}$ be some neighborhood of $x^{*}$. Our goal is to prove $\mathbb{P}(X^{\operatorname{SGD}}_{\lim}\in U)>0$. Choose $\gamma\in\mathbb{R}$, such that $e^{\lambda(x^{*})}<\gamma<1$ and let $\delta>0$ and $C_{\gamma}:\Omega\to\mathbb{R}\cup\infty$ be such that the conclusion of Lemma 3.6 holds, i.e. such that $\mathbb{P}(\mathbb{C}_{\gamma}(\omega)<\infty)=1$ and that we have
@@ 3.7 Generated Matrix Semi-groups
It remains to prove Theorem B (ii). The proof of Theorem B (i) in the previous section was of a quenched nature. One might expect that the best approach to proving Theorem B (ii) is to construct $\omega$-wise center-stable manifolds similar to the ones constructed in section 3.4. While an invariant manifold theory (Pesin theory) has been developed for random dynamical systems (see e.g. Liu and Qian (1995) or Arnold (1998)), these results only provide center-stable manifolds for single points $x\in\mathcal{M}$. For the argument in section 3.4 it was crucial to have a center-stable manifold for an open subset of $\mathcal{M}$. The authors are not aware of any method to construct such a random center-stable manifold.
@@
We will show this by constructing a Lyapunov function (see e.g. Benaïm and Hurth (2022)) defined on a neighborhood of $x^{*}$ which goes to infinity near $\mathcal{M}$. This is inspired by previous work on the instability of invariant subspaces for stochastic differential equations (cf. Baxendale (1991); Baxendale and Stroock (1988); Bedrossian et al. (2022); Blumenthal et al. (2023); Coti Zelati and Hairer (2021)). While these works rely on conditions of Hörmander-type to establish the existence of a spectral gap in the so-called projective process, such a spectral gap is excluded by the discrete nature of our problem. Instead we use an argument due to Le Page (1982) (cf. also Bougerol and Lacroix (1985) for a survey in English) to find a spectral gap in a different topology. Le Page’s argument needs the matrix semi-group on which the linear co-cycle $\Psi$ is supported to satisfy two algebraic properties, namely being contracting and strongly irreducible (cf. Definition 3.8 below). In this section, we will show that they follow from the regularity of $x^{*}$.
@@
Proof  Since $x^{*}\in\mathcal{M}$ is regular, in particular
@@
Thus $\mathcal{S}_{1}(x^{*})\subseteq\operatorname{GL}(N,\mathbb{R})$ and since $\mathcal{S}(x^{*})$ is generated by $\mathcal{S}_{1}(x^{*})$ also $\mathcal{S}(x^{*})\subseteq\operatorname{GL}(N,\mathbb{R})$.
Definition 3.8
@@
Proof  Let $x^{*}\in\mathcal{M}$ be a regular point with $\lambda(x^{*})>0$. For ease of notation, we will write $G:=G_{x^{*}}$ and $G_{[i]}:=G_{x^{*},[i]}$. We start by proving that $\mathcal{S}(x^{*})$ is contracting. From the original definition (2.12) of $\lambda(x^{*})$, we get the inequality
@@
Since $G_{1,1}\neq\frac{2}{\eta}$, we have $|1-\eta G_{1,1}|\neq 1$ and the sequence $(\kappa_{n})$ consists of pairwise distinct elements. Thus, in particular, the sequence $(W_{n})$ consists of pairwise distinct subspaces, completing the proof.
@@ 3.8 Construction of Lyapunov Functions
Following recent works in fluid dynamics (Bedrossian et al. (2022); Blumenthal et al. (2023)), we will study the family of linear operators $(\mathcal{P}_{q}:\mathcal{C}^{0}(S^{N-1})\to\mathcal{C}^{0}(S^{N-1}))_{q\in\mathbb{R}}$ given by
@@
Here $\mathcal{C}^{0}(S^{N-1})$ denotes the Banach space of real-valued continuous functions on $S^{N-1}$, equipped with the supremum-norm. It can be readily seen that the operators $\mathcal{P}_{q}$ are bounded in this norm [^15] In fact $\|\mathcal{P}_{q}\|\leq K_{2}^{q}$ for $q\geq 0$ and $\|\mathcal{P}_{q}\|\leq K_{1}^{-q}$ for $q\leq 0$, with $K_{1}$ and $K_{2}$ defined by (3.28) and (3.29).. The arguments in Bedrossian et al. (2022) and Blumenthal et al. (2023) rely on the fact that $\mathcal{P}_{0}$ admits a spectral gap in their settings. Unfortunately the discrete nature of our setting prohibits such a spectral gap, at least in $\mathcal{C}^{0}(S^{N-1})$. In order to circumvent this obstacle we must consider a different Banach space.
@@
Lemma 3.10 (Proposition V.4.1 in Part A of Bougerol and Lacroix (1985))
@@
For the proof, we refer to Bougerol and Lacroix (1985). The integrability assumption is satisfied trivially in our setting, as $\Psi^{(1)}_{\omega}$ can only take finitely many values.
Lemma 3.10 (ii) says that 1 an dominant eigenvalue of $\mathcal{P}_{0}$, i.e. it is isolated and the spectral value with the largest absolute value. By classical perturbation theory (Kato (1995)) this implies that for $q$ sufficiently close to 0 the largest spectral value of $\mathcal{P}_{q}$ is also an isolated eigenvalue and both the dominant eigenvalue and the corresponding Riesz projections are analytic in $q$. Put more precisely, we get the following corollary (cf. also Theorem V.4.3 in Part A of Bougerol and Lacroix (1985)). Here $\dot{\mathcal{C}}^{\alpha}$ denotes the dual space of $\mathcal{C}^{\alpha}$ and $\mathcal{Q}^{*}\in L(\dot{\mathcal{C}}^{\alpha})$ the dual operator of an operator $\mathcal{Q}\in L(\mathcal{C}^{\alpha})$.
Corollary 3.11
@@
The arguments for the rest of the section are similar to the ones made in chapter 4 of Bedrossian et al. (2022). Henceforth, let $\alpha\in(0,1)$ be as in Lemma 3.10.
Theorem 3.12
@@
Proof  Note that (3.42) states that $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$. In the following, we will show that $\frac{\mathrm{d}}{\mathrm{d}q}r(q)|_{q=0}=\lambda(x^{*})>0$. This will imply that for sufficiently small $p>0$, we can set $\gamma:=r(-p)<1$ and $f^{*}:=f_{-p}$. Since, for small $p$, the function $f_{-p}$ is close161616in the $\mathcal{C}^{\alpha}$-sense, but thus in particular in the $\mathcal{C}^{0}$-sense to $f_{0}=\mathbf{1}$, it is indeed a positive function and by (3.41), we have $\mathcal{P}_{-p}f^{*}=\gamma f^{*}$.
@@
By the Oseledec theorem (cf. Theorem 2.1) the limit in the last line is equal to $\lambda(x^{*})$. Thus we get $r(q)\geq e^{q\lambda(x^{*})}$, for sufficiently small $q$. Since we already know that $r$ is differentiable in 0 from Corollary 3.11 and since $r(0)=1$, this implies $\frac{\mathrm{d}}{\mathrm{d}q}r(q)|_{q=0}=\lambda(x^{*})$, completing the proof.
@@
Proof  This is a direct consequence of (3.42). Let $w\in\mathbb{R}^{N}\setminus\{0\}$ and set $s:=\frac{w}{\|w\|}\in S^{N-1}$. Then
@@
Proof  Using $\alpha$-Hölder continuity of $f^{*}$ and the bounds in (3.45), we can estimate
@@
Proof  Let $\varepsilon>0$ and choose $\tilde{\delta}>0$ such that the conclusion (3.47) of the previous lemma holds. Recall that (cf. (3.28))
@@
which implies (3.49).
Corollary 3.16
@@
Proof  First note that if we choose $\delta<K_{1}$, using Corollary 3.5, we get
@@
for all $\omega\in\Omega$ and all ${(v,w)\in V_{\delta}\setminus(\mathbb{R}^{D-N}\times\{0\})}$.